Calculate the Pearsons correlation coefficient
1
Expert's answer
2019-05-02T10:02:10-0400

Pearson's correlation coefficient is the covariance of the two variables divided by the product of their standard deviations. 


"\rho (X, Y) = \frac{cov(X, Y)} {\sigma(X) ×\sigma(Y) } ,""\sigma(X) = \sqrt{E[(X - E[X])^2]} ,"


"\sigma(Y) = \sqrt{E[(Y - E[Y])^2]} ,""cov (X, Y) = E[XY] - E[X]\u00d7E[Y],"

where E is mean.